Circular No. 8585

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005dj IN UGC 3545 T. Boles, Coddenham, England, reports the discovery of an apparent

supernova (mag 16.6) on unfiltered CCD images taken on Aug. 18.086 and 19.001 UT with a 0.35-m reflector. The new object is located at α = 6^{h}49^{m}58^{s}.41, δ = +61^{o}37′08′′.5 (equinox 2000.0), which is ≈ 20′′.0 east and 16′′.5 south of the center of UGC 3545. Nothing is visible at this location on Boles’ images from 2005 Apr. 3 and Jan. 2 (limiting mag 19.5) or on Digitized Sky Survey images from 1998 (limiting red mag 21.0) and 1995 (limiting blue mag 20.0).

SUPERNOVAE 2005dg AND 2005dh M. Salvo, B. Schmidt, T. Davis, and S. Sankarankutty, Australian Na-

tional University (ANU), report that a spectrogram (range 350–900 nm), obtained with the ANU 2.3-m telescope (+ Double-Beam Spectrograph) on Aug. 15.40 UT, shows SN 2005dh (IAUC 8583) to be a type-Ia supernova near maximum light; adopting the NED recession velocity for the host galaxy (from the RC3 Catalogue) the expansion velocity of the Si II 635.5-nm line is around 16600 km/s. A spectrogram (range 340–900 nm), obtained on Aug. 16.70, shows SN 2005dg (IAUC 8581) to be a type-Ic supernova a few days past maximum light.

COMETS C/2005 O4, O5, O6, P1, P2 (SOHO) K. Battams (Interferometrics, Inc., and Naval Research Laboratory)

reports measurements for additional tailless comets found on SOHO website images (cf. IAUC 8584); all were Kreutz sungrazers except for C/2005 O5 (Meyer group). All five objects had a stellar appearance except for C/2005 P1 (which appeared very diffuse only in the single C2 image, but stellar in the C3 images, where it reached mag ∼ 6.1 at 12.0R_{⊙} on Aug. 6.071 UT). C/2005 O5 and C/2005 P2 were extremely faint. C/2005 O4 reached mag ∼ 6.5 at 12.3R_{⊙} on July 24.512. C/2005 O6 reached mag ∼ 5.5 at 10.6R_{⊙} on July 26.888.

Comet 2005 UT α_{2000} δ_{2000} Inst. F MPEC
C/2005 O4 July 23.971 7^{h}54.^{m}6 +17^{o}35′ C3  TH 2005-Q04
C/2005 O5 26.079 8 14.6 +20 37 C2 TH 2005-Q04
C/2005 O6  26.096 8 01.3 +16 43 C3/2 TH 2005-Q04
C/2005 P1 Aug. 5.431 8 39.5 +14 57 C3/2 TS 2005-Q04
C/2005 P2  5.431 8 44.3 +15 32 C3  TS 2005-Q04
2005 August 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green